# Timer-XL: Long-Context Transformers for Unified Time Series Forecasting

### v2

@@ Abstract
We present Timer-XL, a generative Transformer for unified time series forecasting. To uniformly predict 1D and 2D time series, we generalize next token prediction, predominantly adopted for causal generation of 1D sequences, to multivariate next token prediction. The proposed paradigm uniformly formulates various forecasting scenarios as a long-context generation problem. We opt for the generative Transformer, which can capture global-range and causal dependencies while providing contextual flexibility, to implement unified forecasting on univariate series characterized by non-stationarity, multivariate time series with complicated dynamics and correlations, and covariate-informed contexts that include both endogenous and exogenous variables. Technically, we propose a universal TimeAttention to facilitate generative Transformers on time series, which can effectively capture fine-grained intra- and inter-series dependencies of flattened time series tokens (patches) and is further strengthened by position embeddings in both temporal and variable dimensions. Timer-XL achieves state-of-the-art performance across challenging forecasting benchmarks through a unified approach. As a large time series model, it demonstrates notable model transferability by large-scale pre-training, as well as contextual flexibility in token lengths, positioning it as a one-for-all forecaster.
@@ 1 Introduction
Transformers have contributed significantly to the fields of natural language and computer vision (Radford et al., 2018; Dosovitskiy et al., 2020), and been extensively applied in time series forecasting, becoming the foundation of specialized forecasters (Zhou et al., 2021; Wu et al., 2021) and large models (Das et al., 2023). As a typical generative task, the quality of predictions heavily relies on the context (Dai et al., 2019). Reliable predictions are made by thoroughly considering endogenous temporal variations and retrieving relevant exogenous correlations into the context (Box, 2013). Further, the context length of training, which serves as an indicator of scaling (Kaplan et al., 2020), determines the maximum input and output of generative Transformers, ultimately enabling long-sequence, high-resolution, and high-frequency generation (Yin et al., 2023; Wang et al., 2024a).
@@
However, existing Transformers in the time series field crucially encounter the context bottleneck. As shown in Figure 1, unlike Transformers for natural language and vision that learn dependencies among thousands to millions of tokens (Kirillov et al., 2023; OpenAI, 2023), time-series Transformers typically work around limited contexts of up to hundreds of time series tokens (patches) (Nie et al., 2022). For univariate time series, a short context length leads to an insufficient perception of global tendencies, overlooking widespread non-stationarity in real-world time series (Hyndman, 2018). The excessive reliance on stationarization, such as normalization (Kim et al., 2021), restricts the model capacity and leads to overfitting of Transformers (Liu et al., 2022b). Moreover, instead of regarding multivariate time series as independent channels (Nie et al., 2022), increasing Transformers explicitly capture intra- and inter-channel dependencies (Zhang & Yan, 2022; Liu et al., 2023; 2024a), leading to an urgency to increase the context length to encompass inter-correlated variables.
@@
Recently, generative Transformers, which present a predominant scalable choice of large language models (Zhao et al., 2023) characterized by the decoder-only architecture, have gained increasing attention in the development of large time series models (Rasul et al., 2023; Ansari et al., 2024) due to their generalization performance and contextual flexibility, that is, one Transformer accommodates all input lengths during inference (Liu et al., 2024b). Therefore, training on longer contexts not only empowers them with the fundamental capability to incorporate more contextual information but also enhances the model versatility toward a one-for-all foundation model, which regards any-variate and any-length time series as one context. Even if previous work has achieved unified modeling on flattened tokens in encoder-only Transformers (Woo et al., 2024), the full dependencies among tokens are incompatible with the causality of forecasting, supported by our experiments revealing that their supervised and generalization performance is still inferior to those of decoder-only ones.
@@
In pursuit of unified time series forecasting, we propose multivariate next token prediction. It unifies forecasting tasks of Figure 1 into the patch-level generation based on long-context sequences. To fully leverage the global-range modeling ability and contextual flexibility of generative Transformers, we develop TimeAttention that learns causal temporal dependencies under channel dependence, elegantly implemented as a masking mechanism without modifying self-attention. With incorporated relative position embeddings for multivariate series, TimeAttention is further enhanced with chronological awareness of time points and the distinction of variables. We propose Timer-XL as the extra long version of a generative time-series Transformer (Timer) (Liu et al., 2024c). We firstly enlarge the context to thousands of tokens and achieve state-of-the-art in univariate, multivariate, and covariate-informed forecasting. With notable contextual flexibility, Timer-XL is generalizable across different temporal dynamics, variables, and datasets. Our contributions lie in three aspects:
@@
  • We propose multivariate next token prediction and unified time series forecasting, strengthening Transformers on enriched forecasting contexts to make reliable predictions.
@@
  • We introduce TimeAttention, a novel causal self-attention tailored for the time series modality on our proposed paradigm, which enables intra- and inter-series modeling with position perception, and maintains the causality and flexibility of generative Transformers.
  • We present Timer-XL that achieves state-of-the-art in well-acknowledged benchmarks. We establish new benchmarks, conduct large-scale pre-training, and provide extensive analyses, demonstrating that long-context Transformers are one-for-all forecasters.
@@ 2 Related Work
Based on the global-range modeling ability of tokens, Transformers have shown great power in time series, rapidly standing out from RNN-, CNN- and MLP-forecasters, especially on long sequences. Initial Transformer-based forecasters primarily focused on long-term prediction, aiming to extend the forecasting horizon while mitigating the quadratic computational growth associated with increasing sequence length (Li et al., 2019; Zhou et al., 2021; Wu et al., 2021). However, the context (lookback) length of previous models is not growing in pace, which hinders Transformers from making fully informed predictions, producing oversmooth results (Liu et al., 2022b). Meanwhile, another advancement has occurred from univariate to multivariate forecasting. Unlike natural language, time series are high-dimensional and inherently correlated (Hyndman, 2018), further requiring large contexts that contain relevant endogenous and exogenous variables. To effectively exploit the intra- and inter-series dependencies, tokenization of Transformers has been extensively developed in temporal-wise (Lim et al., 2021), patch-wise (Nie et al., 2022), and variable-wise (Liu et al., 2023) approaches, with deftly designed Transformers for inter-series modeling (Zhang & Yan, 2022; Wang et al., 2024b).
@@
Despite the prevalent issue of insufficient context in the current situation, few works highlight that these challenges can be uniformly tackled by long-context Transformers. Consequently, we leverage the fundamental sequence modeling capability of the vanilla Transformer, unify various forecasting tasks into multivariate next token prediction, and significantly extend the context length.
@@
Besides, time-series Transformers have experienced an evolution from small scenario-specific models to large foundation models (Das et al., 2023; Woo et al., 2024; Ansari et al., 2024). Among them, decoder-only Transformers, predominantly adopted as the backbone of large language models (Touvron et al., 2023; OpenAI, 2023), have positioned as a scalable and generalizable choice for general time series analysis (Liu et al., 2024c). By independently predicting each token based on preceding contexts, decoder-only models are also multi-length forecasters (Liu et al., 2024b), avoiding resource-intensive training and storage. However, existing decoder-only forecasters are typically trained on a single channel (Nie et al., 2022), making them inaccessible to inter-series dependencies.
@@
Prior work has employed encoder-only Transformers to fully capture dependencies within 2D time series (Woo et al., 2024). The incompatibility between this architecture and next token prediction constrains the flexibility and performance of Transformers. To date, the implementation of next token prediction and multivariate time series forecasting in a single Transformer remains a fundamental challenge. We cope with this issue by incorporating the variable dependency and temporal causality into a fine-grained attention mask, thereby leveraging contextual flexibility and capturing inter-series dependencies simultaneously without altering the self-attention mechanism.
@@ 3.1 Timer
For simplicity, we omit the block index $l$. Timer adopts $\mathbf{W}_{e},\mathbf{W}_{d}\in\mathbb{R}^{D\times P}$ that independently embed and project the token embeddings as $\mathbf{H}=\{\mathbf{h}_{i}\}\in\mathbb{R}^{T\times D}$. $\operatorname{TrmBlock}$ includes feed-forward network and self-attention with the temporal causal mask $\mathcal{T}\in\mathbb{R}^{T\times T}$. $\mathbf{h}_{i}\in\mathbb{R}^{D}$ is the contextual representation of the previous $i$ tokens. Predicted $\hat{\mathbf{x}}_{i+1}$ are supervised with ground truth by MSE loss.
@@ 3.2 Generalize 1D Sequences to 2D Time Series
Compared with Equation 2, the multivariate context length increases from $T$ to $NT$. By contrast, the benefit is that this paradigm learns causal dependencies within each sequence while incorporating fine-grained variable correlations from other sequences, making it a universal forecasting mechanism and outperform channel independence and coarse-grained variable-wise modeling experimentally.
@@
Position embedding has not been sufficiently explored in time-series Transformers. To address the permutation-invariance inherent in Transformers, the attention mechanism should reflect the relative position (chronological information) of tokens that vary on the temporal dimension. Conversely, shuffling the input order of variables should not affect anything other than the output order of variables, necessitating that variate embedding should be permutation-invariant.
To meet the above requirements, we adopt RoPE (Su et al., 2024), a widely utilized relative position embedding in LLMs, as the temporal embedding. In terms of the variable dimension, we use two learnable scalars to keep the reflexivity of variables (Woo et al., 2024). Beyond simply incorporating them together, we provide detailed ablations in Section 4.5 to demonstrate the effectiveness:
@@
Figure 2: Illustration of TimeAttention. For univariate series, temporal mask $\mathcal{T}$ keeps the causality. Given multivariate time series tokenized in a temporal-first order, we adopt the variable dependency $\mathcal{C}$, an all-one matrix, as the left-operand of Kronecker product, expanding temporal mask to a block matrix, which exactly reflects dependencies of multivariate next token prediction. The formulation is also generalizable to univariate and covariate-informed contexts with pre-defined variable dependency.
@@
In contrast to variable-wise (Liu et al., 2023) and non-causal patch-wise tokens (Nie et al., 2022; Woo et al., 2024), our TimeAttention can capture causal patch-wise dependencies within and among all variables. We mask out non-causal dependencies in the calculated attention map $\mathcal{A}$. Before masking, we reorder the tokens by flattening their 2D indices in the temporal-first manner, which is illustrated in the upper left of Figure 2.
We provide an intuitive example to illustrate the causal dependencies within multivariate time series: considering the 2nd token of time series A. To predict its next token, its representation $\mathbf{h}$ should be exactly dependent on the tokens-$\{1,2,4,5\}$. Similarly, we provide the causal dependencies of each token in the upper middle of Figure 2. Starting from the observation of coarse-grained variable dependency, where all variables are inter-correlated in this multivariate task, we find the fine-grained masking on $\mathcal{A}$ can be formally derived as the Kronecker product of two matrices: the adjacency matrix of the variable dependency graph as $\mathcal{C}\in\mathbb{R}^{N\times N}$, and the causal temporal mask $\mathcal{T}\in\mathbb{R}^{T\times T}$:
@@
In multivariate forecasting, the variable dependency forms the complete graph, presenting an all-one matrix $\mathcal{C}$. By generalizing TimeAttention on multiple sequences, generative Transformers can leverage contextual flexibility to encompass relevant covariates as well. In this case, Timer-XL is adapted in two steps: (1) formulate the customized variable dependency as $\mathcal{C}$ and (2) optimize the model using the supervision of target variables. An example (target-$A$-covariate-$B$) of attention is illustrated on the right of Figure 2. In a nutshell, we introduce position embedding and extend the context length of generative Transformers. Toward unified time series forecasting, we flatten 2D time series into a long context and capture fine-grained causal token dependencies.
@@ 4 Experiments
We conduct thorough evaluations of the performance and generality of Timer-XL. Given the saturated performance on previous becnchmarks (Wu et al., 2021; Nie et al., 2022; Liu et al., 2023), where Transformers tend to overfit on these tasks due to their data-hungry nature (Touvron et al., 2021), we establish new long-context benchmarks and will release our datasets for the advancement of this field. Detailed datasets, baseline models, and experimental configurations are provided in Appendix B. We also dive into commonly adopted techniques, such as channel independence (Nie et al., 2022) and normalization (Kim et al., 2021). We conclude that long-context Transformers can tackle these challenges without any task-specific designs.
@@ 4.1 Univariate Time Series Forecasting
Due to the insufficient dataset length when extending contexts in univariate datasets (Makridakis et al., 2020), we adopt wide-recognized benchmarks from Liu et al. (2023). Although these datasets are originally multivariate, they will be predicted in a univariate approach with the implementation of channel independence. Different from the previous long-term forecasting setting, we focus on reliable prediction based on a long context. Thus, we fix the prediction horizon and increase the lookback length to monthly and yearly levels. We also establish a long-context univariate benchmark based on the challenging 40-year ECMWF Reanalysis v5 dataset (Hersbach et al., 2020), where yearly contexts are adopted to predict the land-surface temperature of a single site (ERA5-S).

### v3

@@ Abstract
We present Timer-XL, a generative Transformer for unified time series forecasting. To uniformly predict 1D and 2D time series, we generalize next token prediction, predominantly adopted for causal generation of 1D sequences, to multivariate next token prediction. The proposed paradigm uniformly formulates various forecasting scenarios as a long-context generation problem. We opt for the generative Transformer, which can capture global-range and causal dependencies while providing contextual flexibility, to implement unified forecasting on univariate series characterized by non-stationarity, multivariate time series with complicated dynamics and correlations, and covariate-informed contexts that include both endogenous and exogenous time series. Technically, we propose a universal TimeAttention to facilitate generative Transformers on multiple time series, which can effectively capture fine-grained intra- and inter-series dependencies of flattened time series tokens (patches), and is further enhanced by deftly designed position embeddings for the temporal and variable dimensions. Timer-XL achieves state-of-the-art performance across challenging forecasting benchmarks through a unified approach. Based on large-scale pre-training, Timer-XL also demonstrates notable zero-shot performance, making it a promising architecture for large time series models.
@@ 1 Introduction
Transformers have contributed significantly to the fields of natural language and computer vision (Radford et al., 2018; Dosovitskiy et al., 2020), and been extensively applied in time series forecasting, becoming the foundation of specialized forecasters (Zhou et al., 2021; Wu et al., 2021) and large models (Das et al., 2023). As a typical generative task, the quality of predictions heavily relies on the context (Dai et al., 2019). Reliable predictions are made by thoroughly considering endogenous temporal variations and retrieving relevant exogenous correlations into the context (Box, 2013). Further, the pre-training context length, which serves as an indicator of scaling (Kaplan et al., 2020), determines the maximum input and output of generative Transformers, ultimately enabling long-sequence, high-resolution, and high-frequency generation (Yin et al., 2023; Wang et al., 2024a).
@@
However, existing Transformers in the time series field crucially encounter the context bottleneck. As shown in Figure 1, unlike Transformers for natural language and vision that learn dependencies among thousands to millions of tokens (Kirillov et al., 2023; OpenAI, 2023), time-series Transformers typically work around limited contexts of up to hundreds of time series tokens (patches) (Nie et al., 2022). For univariate time series, a short context length leads to an insufficient perception of global tendencies, overlooking widespread non-stationarity in real-world time series (Hyndman, 2018). The excessive reliance on stationarization, such as normalization (Kim et al., 2021), restricts the model capacity and leads to overfitting of Transformers (Liu et al., 2022b). Moreover, instead of regarding multivariate time series as independent channels (Nie et al., 2022), increasing Transformers explicitly capture intra- and inter-channel dependencies (Zhang & Yan, 2022; Liu et al., 2023; 2024a), leading to an urgency of extending the context length to encompass inter-correlated variables.
@@
Recently, generative Transformers, which present a predominant scalable choice of large language models (Zhao et al., 2023) characterized by the decoder-only architecture, have gained increasing attention in the development of large time series models (Rasul et al., 2023; Ansari et al., 2024). Based on their generalization performance and contextual flexibility, one model can accommodate varing input lengths during inference (Liu et al., 2024b). Therefore, pre-training on longer contexts not only empowers them with the fundamental capability to incorporate more contextual information but also enhances the model versatility toward a one-for-all foundation model, which regards any-variate and any-length time series as one context. Even if previous work (Liu et al., 2024a) has achieved unified modeling on flattened tokens in encoder-only Transformers (Nie et al., 2022), our empirical results in Figure 3 reveal that encoder-only Transformers encounter performance bottleneck especially on long-context time series, while generative Transformers can mitigate performance degradation well.
@@
In pursuit of unified time series forecasting, we propose multivariate next token prediction. It unifies forecasting tasks of Figure 1 into the patch-level generation based on long-context sequences. To fully leverage the global-range modeling ability and length-flexibility of generative Transformers, we develop TimeAttention that captures causal temporal dependencies in the channel-dependent approach, presenting a novel masking mechanism from the disentangling perspective. With incorporated relative position embeddings for multivariate series, TimeAttention is aware of the chronological order of time points and achieves permutation-equivariance (Zaheer et al., 2017) of variables. We propose Timer-XL as the extra long version of a generative time-series Transformer (Timer) (Liu et al., 2024c). We enlarge the context to thousands of patch tokens and achieve state-of-the-art in univariate, multivariate, and covariate-informed forecasting. With comparable pre-training scale and model size, Timer-XL outperforms advanced large models in zero-shot forecasting. Our contributions lie in three aspects:
@@
  • We propose multivariate next token prediction and unified time series forecasting, strengthening Transformers with enriched forecasting contexts to make reliable predictions.
@@
  • We introduce TimeAttention, a novel causal self-attention tailored for the time series modality on our proposed paradigm, which enables intra- and inter-series modeling with positional awareness, and maintains the causality and length-flexibility of generative Transformers.
  • We propose Timer-XL, a versatile Transformer for one-for-all forecasting, which mitigates performance degradation in long-context time series, achieves state-of-the-art performance in task-specific benchmarks, and presents notable zero-shot results by pre-training.
@@ 2 Related Work
Based on the global-range modeling ability of tokens, Transformers have shown great power in time series, rapidly standing out from RNN-, CNN- and MLP-forecasters, especially on long sequences. Initial Transformer-based forecasters primarily focused on long-term prediction, aiming to extend the forecasting horizon while mitigating the quadratic computational growth associated with increasing sequence length (Li et al., 2019; Zhou et al., 2021; Wu et al., 2021). However, the context (lookback) length of previous models is not growing in pace, which hinders Transformers from making fully informed predictions, producing oversmooth results (Liu et al., 2022b). Meanwhile, another advancement has emerged from univariate to multivariate forecasting. Unlike natural language, time series are high-dimensional and inherently correlated (Hyndman, 2018), further requiring longer contexts that contain relevant endogenous and exogenous variables. To effectively exploit the intra- and inter-series dependencies, tokenization of Transformers has been extensively developed in temporal-wise (Lim et al., 2021), patch-wise (Nie et al., 2022), and variable-wise (Liu et al., 2023) approaches, with deftly designed Transformers for inter-series modeling (Zhang & Yan, 2022; Wang et al., 2024b).
@@
Despite the challenges caused by the limited context for making predictions, few works highlight that these challenges can be uniformly tackled by long-context generative Transformers. Consequently, we leverage the fundamental sequence modeling capability of the vanilla Transformer, extend the context length significantly, and unify various forecasting tasks into multivariate next token prediction.
@@
Recently, time-series Transformers have experienced the evolution from small task-specific models to large foundation models (Das et al., 2023; Woo et al., 2024; Ansari et al., 2024). Among them, decoder-only Transformers, predominantly adopted as the backbone of large language models (Touvron et al., 2023; OpenAI, 2023), have positioned as a scalable and generalizable choice for general time series analysis (Liu et al., 2024c). By independently predicting each token based on preceding contexts, decoder-only models are also multi-length forecasters (Liu et al., 2024b), avoiding resource-intensive training and lookback-search. However, existing decoder-only forecasters are generally trained on a single channel (Nie et al., 2022), making them inaccessible to inter-series dependencies.
@@
Prior work has employed encoder-only Transformers to fully capture dependencies within 2D time series (Liu et al., 2024a). The incompatibility between this architecture and temporal causality may constrain the flexibility and performance of Transformers. To date, the implementation of next token prediction and multivariate time series forecasting in a single Transformer remains a fundamental challenge. Our work addresses this issue by disentangling the fine-grained token dependencies clearly into variable dependencies and temporal causal masks, thereby maintaining temporal causality and capturing inter-series dependencies simultaneously without altering the self-attention mechanism.
@@ 3.1 Timer
For simplicity, we omit the block index $l$. Timer adopts $\mathbf{W}_{e},\mathbf{W}_{d}\in\mathbb{R}^{D\times P}$ that independently embed and project the token embeddings as $\mathbf{H}=\{\mathbf{h}_{i}\}\in\mathbb{R}^{T\times D}$. $\operatorname{TrmBlock}$ includes feed-forward network and self-attention with the temporal causal mask $\mathcal{T}\in\mathbb{R}^{T\times T}$. $\mathbf{h}_{i}\in\mathbb{R}^{D}$ is the contextual representation of the previous $i$ tokens. Predicted $\hat{\mathbf{x}}_{i+1}$ are supervised by ground truth in MSE loss.
@@ 3.2 Generalize 1D Sequences to 2D Time Series
Compared with Equation 2, the multivariate context length increases from $T$ to $NT$. By contrast, the benefit is that this paradigm learns causal dependencies within each sequence while incorporating exogenous variable correlations from other sequences, making it a universal forecasting paradigm that outperforms channel independence or coarse-grained variable-wise modeling experimentally.
@@
Position embedding has not been sufficiently explored in time-series Transformers. To avoid the permutation-invariance of self-attention, positional embedding is required to reflect the chronological order of tokens on the temporal dimension. As for the variable dimension, shuffling the input order of variables should not affect anything other than the output order of variables. Formally, the processing on multiple variables should be permutation-equivalent (Zaheer et al., 2017).
To meet the above requirements, we adopt RoPE (Su et al., 2024), a widely utilized position embedding on the temporal dimension. For the variable dimension, we use two learnable scalars in each head to keep the permutation-equivalence of variables (Woo et al., 2024). Beyond simply incorporating them together, we provide detailed ablations in Section E.3 to demonstrate the effectiveness:
@@
Figure 2: Illustration of TimeAttention. For univariate series, temporal mask $\mathcal{T}$ keeps the causality. Given multivariate patch tokens sorted in a temporal-first order, we adopt the variable dependencies $\mathcal{C}$, an all-one matrix, as the left-operand of Kronecker product, expanding temporal mask to a block matrix, which exactly reflects dependencies of multivariate next token prediction. The formulation is also generalizable to univariate and covariate-informed contexts with pre-defined variable dependency.
@@
In contrast to variable-wise (Liu et al., 2023) and non-causal patch-wise tokens (Nie et al., 2022; Woo et al., 2024), our TimeAttention aims to capture causal patch-wise dependencies within and among all variables. Concretely, we sort patch tokens by flattening their 2D indices into 1D indices in the temporal-first manner, which is illustrated in the upper left of Figure 2. Note that the order of variables does not matter, since Equation 6 guarantees their permutation-equivalence.
We provide an intuitive example to illustrate the causal dependencies within multivariate time series: considering the 2nd token of time series A. To predict its next token, its representation $\mathbf{h}$ should be exactly dependent on the tokens-$\{1,2,4,5\}$. Similarly, we provide all causal dependencies of each token in Figure 12. Based on the visualized attention mask and variable dependencies presented in Figure 2, where all variables are inter-correlated, we find that causal temporal dependencies in $\mathcal{A}$ can be formally disentangled by the Kronecker product into (1) the adjacency matrix of the variable dependency graph $\mathcal{C}\in\mathbb{R}^{N\times N}$ and (2) the causal temporal mask $\mathcal{T}\in\mathbb{R}^{T\times T}$:
@@
In multivariate forecasting, the variable dependency forms the complete graph, presenting an all-one matrix $\mathcal{C}$. By generalizing TimeAttention on multiple sequences, generative Transformers can leverage its length-flexibility to encompass relevant covariates as well. In this case, Timer-XL is adapted in two steps: (1) formulate the customized variable dependency as $\mathcal{C}$ and (2) optimize the model using the supervision of target variables. An example (target-$A$-covariate-$B$) of TimeAttention is illustrated on the right of Figure 2. In a nutshell, we adopt position embeddings for the temporal and variable dimensions. To achieve unified time series forecasting, we flatten 2D time series into a unified context and capture fine-grained causal token dependencies.
@@ 4 Experiments
We conduct thorough evaluations of the performance and versatility of Timer-XL. Given that the long-context forecasting paradigm receives less attention in the community, which can be concealed due to the performance saturation on previous benchmarks (Wu et al., 2021; Nie et al., 2022), we delve into the backbone and representation comparison on Transformers. New long-context benchmarks are established, which will be released for the advancement of this field. Detailed datasets, baseline models, and experimental configurations are provided in Appendix B. We also dive into commonly adopted techniques, such as channel independence (Nie et al., 2022) and normalization (Kim et al., 2021). We conclude that generative Transformers can tackle these challenges without specific designs.
@@ 4.1 Univariate Time Series Forecasting
Due to the insufficient dataset length when extending contexts in univariate datasets (Makridakis et al., 2020), we adopt wide-recognized benchmarks from Liu et al. (2023). Although these datasets are originally multivariate, they aim to be predicted in a univariate approach with the implementation of channel independence. Different from the previous long-term forecasting setting, we focus on reliable prediction based on a long context. Therefore, we fix the prediction horizon and increase the lookback length to monthly and yearly levels. We also establish a long-context univariate benchmark based on the challenging 40-year ECMWF Reanalysis v5 dataset (Hersbach et al., 2020), where yearly contexts are adopted to predict the land-surface temperature of a single site (ERA5-S).